# Dynamic similarity of vortex shedding in a superfluid flowing past a penetrable obstacle

### v1

@@ Abstract
We numerically investigate wake dynamics in a superfluid flowing past a penetrable obstacle. The penetrable obstacle does not fully deplete the density, and we define an effective diameter $D_{\rm eff}$ from the Mach-1 contour of the time-averaged irrotational flow around the obstacle, which delineates the local supersonic region where quantized vortices nucleate. Using this flow-defined length, we construct a superfluid Reynolds number $\mathrm{Re}_{\rm s}\equiv(v_{0}-v_{c})D_{\rm eff}/(\hbar/m)$ with $v_{0}$ being the flow speed, $v_{c}$ as the critical velocity and $m$ as the particle mass, and show that $\mathrm{Re}_{\rm s}$ organizes the wake dynamics across obstacle sizes and strengths: the transition from dipole-row emission to alternating vortex cluster shedding occurs for $\mathrm{Re}_{\rm s}\approx 2$ and the Strouhal number and drag coefficient collapse onto universal curves versus $\mathrm{Re}_{\rm s}$. These results extend the notion of dynamic similarity in superfluid flows to penetrable obstacles and demonstrate that the dynamically relevant length scale is determined by the supersonic region rather than by the geometric obstacle size.
@@ I Introduction
Dynamic similarity refers to the principle that flows, with different absolute scales of length, time, and mass in fluid properties, can exhibit identical behavior when expressed in terms of appropriate dimensionless parameters [1]. A paradigmatic example is a flow past bluff bodies, where wake transitions and forces are oranized according to the Reynolds number $\mathrm{Re}=UD/\nu$ [2], where $U$ is the flow velocity, $D$ is the characteristic length of the body and $\nu$ is the kinematic viscosity. Above $\mathrm{Re}\approx 50$ [3], a periodic alternating vortex pattern, known as the von Kármán vortex street, starts to form, and for larger values of $\mathrm{Re}$, the wake gradually loses strict periodicity, entering turbulent vortex shedding around $\mathrm{Re}=250$ [4]. Nevertheless, the lift force, normal to the direction of the background flow, still has a dominant frequency. The shedding frequency $f_{s}$ is nondimensionalized through the Strouhal number $\mathrm{St}=f_{\mathrm{s}}D/U$, showing universal $\mathrm{St}$$\mathrm{Re}$ trends over broad ranges of $\mathrm{Re}$, where $\mathrm{St}\simeq 0.2$, nearly up to nearly $\mathrm{Re}\sim 10^{5}$ [2, 5]. Moreover, when the drag force $F_{\mathrm{D}}$ is normalized to the coefficient $C_{\mathrm{D}}=F_{\mathrm{D}}/(\tfrac{1}{2}\rho U^{2}A)$, where $\rho$ is the mass density of the fluid and $A$ is the reference area of the object, it exhibits well–known dependence on the Reynolds number [6, 7], typically taking values of $C_{\mathrm{D}}\sim 1$ over a broad range of $\mathrm{Re}\sim 10^{3}-10^{5}$ [4].
@@
The advent of quantum fluids has enabled tests of such universality of dynamic similarity in new regimes. Liquid helium systems have long served as a platform for research on inviscid flows [8, 9, 10], and ultracold atomic gases now provide an especially clean and highly controllable system, where versatile flow geometries can be engineered by optical methods with high precision. In these systems, one can investigate the properties of dissipationless fluids and probe exotic phenomena unique to superfluids, such as the nucleation and dynamics of quantized vortices [11, 12, 13, 14, 15, 16]. An increasing number of studies have highlighted ensuing parallels between classical and quantum flows. The von Kármán vortex streets behind obstacles [17, 18, 19, 20, 21], the scaling law in quantum turbulence [22, 23, 24] and universal scaling along with the superfluid Reynolds number in shedding dynamics [25, 26, 27] represent cases in which well–known concepts in classical fluid dynamics can be extended to quantum cases. Recently, it was theoretically shown that dissipation in classical fluids can be understood through the Josephson–Anderson relation for superfluids [28] presenting a case where knowledge pertaining to superfluids offers valuable insight into classical fluid dynamics.
@@
In this work, we investigate dynamic similarity in the wake dynamics in a superfluid, focusing on a regime involving a penetrable obstacle. The feature by which the fluid can penetrate the body of the obstacle fundamentally modifies the vortex nucleation process [20, 29]. In contrast to impenetrable hard obstacles, which nucleate individual vortices from the boundary of density-depleted regions, penetrable obstacles produce vortex dipoles through a precursor rarefaction pulse. This penetrable obstacle constitutes a regime without a true classical analogue, but it is readily realized in atomic Bose-Einstein condensate (BEC) experiments using a focused laser beam whose potential strength imparted onto the atoms is below the chemical potential of the BEC [20, 29, 31, 27, 30]. There have been experimental studies of penetrable obstacles in which the periodic generation of vortex dipoles from a moving obstacle was observed [29] and the quantitative relationship between the dipole generation rate and the obstacle speed and strength was investigated [27]. Because the dynamic similarity in wake dynamics was numerically presented for impenetrable obstacles [25], it is highly desirable to extend this cnocept to penetrable obstacles. However, a unifying quantitative framework for vortex shedding across different regimes is still lacking, particularly because it is difficult to define the length scale of the penetrable obstacke in an unambiguous manner.
@@
Here, we analyze shedding dynamics from a penetrable obstacle by numerically solving Gross–Pitaevskii equations (GPEs) and present an appropriate effective length scale that captures the evolution of the shedding behaviors. We define the effective diameter from the Mach-1 contour of the time-averaged irrotational flow around the penetrable obstacle, which delineates the locally supersonic region where quantized vortices nucleate. We demonstrate that once this length is incorporated into a new superfluid Reynolds number, the resulting dimensionless quantities, in this case the Strouhal number and drag coefficient, exhibit universal behaviors in close analogy with classical fluids. Our findings thus extend the scope of dynamic similarity to penetrable superfluid flows, showing that even in this uniquely quantum regime, the universal language of fluid dynamics persists.
@@ II Method
We consider a situation in which a BEC flows at constant velocity $\bm{v}_{\infty}=-v_{0}\hat{\mathbf{x}}$ through obstacle potential $V(\mathbf{r})$. To simulate the superfluid vortex shedding induced by the obstacle, we numerically solve the two-dimensional GPE for the condensate wave function $\psi(\mathbf{r},t)$, which is given by
@@
The initial trial wavefunction was prepared from the Thomas–Fermi approximation in the presence of the obstacle and relaxed to the true ground state for a stationary BEC by imaginary–time propagation [32]. The real time evolution is calculated using the Fourier pseudo-spectral method [33] with the fourth order Runge-kutta method with $\Delta t=0.01\tau$ and $\tau=\hbar/\mu$. During real-time evolution, the background flow was linearly increased to target velocity $v_{0}$ over a short acceleration interval of $0.1\,\tau$. The simulation domain is $350<x/\xi<150$ and $|y/\xi|<125$ and the spatial resolution is $0.5\xi$ per grid point, where $\xi=\hbar/\sqrt{m\mu}$ is the condensate healing length.
@@
with $\alpha=x,y$, $\gamma_{0}=0.1$, $w_{x}=25\xi$, $w_{y}=20\xi$ and $d=3\xi$. Furthermore, to prevent vortex reentry under the periodic boundary conditions, we use the vortex unwind algorithm to remove the vortex adjacent to the boundary region [35, 25]. Vortex removal is implemented every $5\tau$ during time evolution.
To investigate the steady situation, we examined the vortex shedding over a window of $2000\tau\leq t\leq 10000\tau$, thus excluding the initial acceleration stage. To break symmetry during initial shedding, we added a small amount of complex random noise to the initial field after the imaginary time evolution, with uniformly distributed real and imaginary parts scaled to a relative amplitude of the order of $10^{-4}$.
@@ III.1 Vortex dipole–to–cluster transition
Figure 1: Vortex shedding in a Bose-Einstein condensate flowing past a penetrable obstacle. The condensate flows from right to left (black arrow) at a speed of $v_{0}$. The Gaussian obstacle has strength $V_{0}/\mu=0.9$ and width $\sigma/\xi=20$. (a)–(c) Superfluid vorticity distributions $\omega_{\mathrm{s}}(\mathbf{r})$ for $v_{0}/c_{s0}=0.25$ (a), $0.30$ (b), and $0.40$ (c), where $c_{s0}$ is the speed of sound in the unperturbed condensate. $\omega_{\mathrm{s}}(\mathbf{r})$ is normalized by $\omega_{\mathrm{s0}}=\rho_{0}/\tau$. As $v_{0}$ increases, shedding evolves from periodic vortex-dipole emission to same-signed vortex-cluster shedding. Dashed circles indicate the obstacle’s $1/e^{2}$ radius, centered at $x=0$. (d)–(f) Time traces of drag force $F_{\mathrm{D}}$ (solid blue) and lift force $F_{\mathrm{L}}$ (dotted red) for the same parameters used in (a)–(c). The force is normalized by $F_{\mathrm{\mu}}=\mu/\xi$. With an increase in $v_{0}$, $F_{\mathrm{D}}$ grows and becomes irregular, while $F_{\mathrm{L}}$ increases in terms of its magnitude and oscillatory strength, reflecting the transition from dipole- to cluster-dominated shedding.
@@
providing a convenient measure of vortex generation, annihilation, and propagation around the obstacle. The force by the obstacle is, according to Ehrenfest’s theorem, estimated as
@@
In Fig. 1, we display $\omega_{\mathrm{s}}(\mathbf{r})$ and $\mathbf{F}(t)$ for different flow velocities with $V_{0}/\mu=0.9$ and $\sigma/\xi=20$. Just above the critical velocity $v_{\mathrm{c}}\approx 0.18c_{\rm s0}$, where $c_{\rm s0}=\sqrt{\mu/m}$ is the speed of sound in the unperturbed condensate, vortex dipoles are nucleated within the penetrable obstacle via phase accumulation and phase slip (Fig. 1(a)[20]. These dipoles form nearly parallel rows as they propagate downstream. Because such dipole rows are intrinsically unstable [38], the downstream wake gradually reorganizes into a staggered V–shaped pattern [21]. As the velocity increases, this instability spreads upstream toward the obstacle, resulting in increasingly irregular vortex configurations in the near wake (Fig. 1(b)). At still higher velocities, the nucleation mechanism itself becomes unstable, exhibiting a qualitative change (Fig. 1(c)): rarefaction pulses fragment into multiple vortices, which subsequently interact, annihilate, or produce Jones–Roberts solitons [39]. These interactions lead to the formation of same–sign vortex clusters that persist downstream and may eventually organize into a quasi–von Kármán street [40], despite the fact that the obstacle is penetrable.
@@
The corresponding force response, shown in Figs. 1(d)1(f), reflects the same progression. In the dipole–row regime near $v_{\mathrm{c}}$, the drag force exhibits periodic oscillations with negligible lift, consistent with symmetric dipole nucleation. As the velocity increases, the drag oscillation frequency rises and the signal fluctuates slightly more, in line with the growing instability in the wake. Once vortex clusters appear, the drag force becomes strongly irregular, while the lift force develops pronounced periodic oscillations associated with alternating cluster formation (Fig. 1(f)). It is clearly revealed that the shedding behavior for a penetrable obstacle changes from periodic dipole–row emission to alternating same–sign vortex cluster shedding as the flow velocity increases, which is consistent with earlier numerical studies [41].
@@
We investigate the dipole–to–cluster transition further by examining the velocity dependence of the drag and lift forces. Fig. 2(a) shows the time-averaged drag force $\overline{F_{\mathrm{D}}}$ and the root-mean-square of the lift force, $(\overline{F_{\mathrm{L}}^{2}})^{1/2}$ as functions of $v_{0}$. Here, the time averages are evaluated for $2000\leq t/\tau\leq 10000$ with snapshots sampled every $\Delta t=10\,\tau$. Both the drag and lift exhibit sudden increases at threshold velocity $v_{\mathrm{th}}\simeq 0.33\,c_{\mathrm{s0}}$, which is also associated with the onset of asymmetric vortex cluster shedding. Interestingly, we find that the drag force closely follows the scaling $\overline{F_{\mathrm{D}}}\sim(v_{0}-v_{\mathrm{c}})^{1/2}$ in the low-velocity dipole-emission regime [42], while it evolves to quadratic dependence $\overline{F_{\mathrm{D}}}\sim(v_{0}-v_{\mathrm{c}})^{2}$ in the high-velocity cluster-shedding regime (Fig. 2(a) inset).
@@
The dipole-to-cluster transition is also examined by analyzing the oscillatory behavior of the drag and lift forces through their Fourier spectra. The drag force displays a well-defined shedding frequency up to $v_{\mathrm{th}}$, consistent with the periodic dipole emission, while above $v_{\mathrm{th}}$, the spectrum becomes irregular and no dominant frequency remains. In contrast, the lift force shows a negligible spectral weight below $v_{\mathrm{th}}$ but develops a clear oscillation frequency above $v_{\mathrm{th}}$, reflecting the alternating release of vortex clusters. From the spectral peaks of the Fourier spectra of $F_{\mathrm{D}}(t)$ and $F_{\mathrm{L}}(t)$, we determine their characteristic oscillating frequencies $f_{\mathrm{D}}$ and $f_{\mathrm{L}}$, respectively. The measurement results are shown in Fig. 2(b) for the corresponding shedding regimes across the threshold velocity. In the dipole shedding regime, $f_{\mathrm{D}}$ increases almost linearly with an increase of $v_{0}$ above $v_{c}$, consistent with previous experiments [27, 29] and in the cluster shedding regime, on the other hand, $f_{\mathrm{L}}$ exhibits very weak dependence on $v_{0}$. Upon crossing the threshold velocity $v_{\mathrm{th}}$, the measured frequencies suggest a relationship of $f_{\mathrm{L}}\approx f_{\mathrm{D}}/2$, which is understandable with the fact that a vortex cluster consists of same-sign vortices.

### v2

@@ Abstract
We numerically investigate wake dynamics in a superfluid flowing past a penetrable obstacle. The penetrable obstacle does not fully deplete the density, and we define an effective diameter $D_{\rm eff}$ from the Mach-1 contour of the time-averaged irrotational flow around the obstacle, which delineates the local supersonic region where quantized vortices nucleate. Using this flow-defined length, we construct a superfluid Reynolds number $\mathrm{Re}_{\rm s}\equiv(v_{0}-v_{c})D_{\rm eff}/(\hbar/m)$ with $v_{0}$ being the flow speed, $v_{c}$ as the critical velocity and $m$ as the particle mass, and show that $\mathrm{Re}_{\rm s}$ organizes the wake dynamics across obstacle sizes and strengths: the transition from dipole-row emission to alternating vortex cluster shedding occurs for $\mathrm{Re}_{\rm s}\approx 2$ and the Strouhal number and drag coefficient collapse onto universal curves versus $\mathrm{Re}_{\rm s}$. These results extend the notion of dynamic similarity in superfluid flows to penetrable obstacles and show that the dynamically relevant length scale is determined by the supersonic region rather than by the geometric obstacle size.
@@ I Introduction
Dynamic similarity refers to the principle that flows, with different absolute scales of length, time, and mass in fluid properties, can exhibit identical behavior when expressed in terms of appropriate dimensionless parameters [1]. A paradigmatic example is a flow past bluff bodies, in which wake transitions and forces are organized according to the Reynolds number $\mathrm{Re}=UD/\nu$ [2], where $U$ is the flow velocity, $D$ is the characteristic length of the body and $\nu$ is the kinematic viscosity. Above $\mathrm{Re}\approx 50$ [3], a periodic pattern of alternating vortices, known as the von Kármán vortex street, starts to form, and for larger values of $\mathrm{Re}$, the wake gradually loses strict periodicity, entering turbulent vortex shedding around $\mathrm{Re}=250$ [4]. Nevertheless, the lift force, normal to the direction of the background flow, still has a dominant frequency. The vortex-shedding frequency $f_{s}$ is nondimensionalized through the Strouhal number $\mathrm{St}=f_{\mathrm{s}}D/U$, showing a universal $\mathrm{St}\text{--}\mathrm{Re}$ relation over a broad range of $\mathrm{Re}$, with $\mathrm{St}\simeq 0.2$ maintained up to $\mathrm{Re}\sim 10^{5}$ [2, 5, 6]. Moreover, when the drag force $F_{\mathrm{D}}$ is normalized to the coefficient $C_{\mathrm{D}}=F_{\mathrm{D}}/(\tfrac{1}{2}\rho U^{2}A)$, where $\rho$ is the mass density of the fluid and $A$ is the reference area of the object, it exhibits well–known dependence on the Reynolds number [7], typically taking values of $C_{\mathrm{D}}\sim 1$ over a broad range of $\mathrm{Re}\sim 10^{3}-10^{5}$ [4].
@@
The advent of quantum fluids has enabled tests of such universality of dynamic similarity in new regimes. Liquid helium systems have long served as a platform for research on inviscid flows [8, 9, 10], and ultracold atomic gases now provide an especially clean and highly controllable system, where versatile flow geometries can be engineered by optical methods with high precision. In these systems, one can investigate the properties of dissipationless fluids and probe exotic phenomena unique to superfluids, such as the nucleation and dynamics of quantized vortices [11, 12, 13, 14, 15, 16]. An increasing number of studies have highlighted parallels between classical and quantum flows. The von Kármán vortex streets behind obstacles [17, 18, 19, 20, 21], the scaling law in quantum turbulence [22, 23] and universal scaling along with the superfluid Reynolds number in shedding dynamics [24, 25, 26] represent cases in which well–known concepts in classical fluid dynamics can be extended to quantum cases. Recently, it was theoretically shown that dissipation in classical fluids can be understood through the Josephson–Anderson relation for superfluids [27], demonstrating that concepts pertaining to superfluid physics can provide valuable insight into classical fluid dynamics.
@@
In this work, we investigate dynamic similarity in the wake dynamics in a superfluid, focusing on a regime involving a penetrable obstacle. The ability of the fluid to penetrate the obstacle fundamentally modifies the vortex nucleation process [20, 28]. In contrast to impenetrable obstacles, which nucleate individual vortices from the boundary of density-depleted regions, penetrable obstacles produce vortex dipoles through a precursor rarefaction pulse. This penetrable obstacle has no true classical analogue, but it is readily realized in atomic Bose-Einstein condensate (BEC) experiments using a focused laser beam whose potential strength experienced by the atoms is below the chemical potential of the BEC [20, 28, 26, 29]. There have been experimental studies of penetrable obstacles in which the periodic generation of vortex dipoles from a moving obstacle was observed [26, 28] and the quantitative relationship between the dipole generation rate and the obstacle speed and strength was investigated [26]. Because the dynamic similarity in wake dynamics was numerically presented for impenetrable obstacles [24], it is therefore natural to expect that this concept can be extended to penetrable obstacles. However, a unifying quantitative framework for vortex shedding across different regimes is still lacking, particularly because it is difficult to define the length scale of the penetrable obstacle in an unambiguous manner.
@@
Here, we analyze shedding dynamics from a penetrable obstacle by numerically solving the Gross–Pitaevskii equations (GPEs) and present an appropriate effective length scale that captures the evolution of the shedding behaviors. We define the effective diameter from the Mach-1 contour of the time-averaged irrotational flow around the penetrable obstacle, which delineates the locally supersonic region where quantized vortices nucleate. We show that, once this length is incorporated into a new superfluid Reynolds number, the resulting dimensionless quantities, in this case the Strouhal number and drag coefficient, exhibit universal behaviors in close analogy with classical fluids. Our findings thus extend the scope of dynamic similarity to penetrable superfluid flows, showing that the framework of dynamic similarity extends to this quantum regime.
@@ II Method
We consider a situation in which a BEC flows at constant velocity $\bm{v}_{\infty}=-v_{0}\hat{\mathbf{x}}$ past an obstacle potential $V(\mathbf{r})$. To simulate the superfluid vortex shedding induced by the obstacle, we numerically solve the two-dimensional GPE for the condensate wave function $\psi(\mathbf{r},t)$, which is given by
@@
The initial trial wavefunction was prepared from the Thomas–Fermi approximation in the presence of the obstacle and relaxed to the true ground state for a stationary BEC by imaginary-time propagation [30]. The real-time evolution is calculated using the Fourier pseudo-spectral method [31] with the fourth-order Runge-kutta method with $\Delta t=0.01\tau$ and $\tau=\hbar/\mu$. During the real-time evolution, the background flow was linearly increased to target velocity $v_{0}$ over a short acceleration interval of $0.1\,\tau$. The simulation domain is $350<x/\xi<150$ and $|y/\xi|<125$ and the spatial resolution is $0.5\xi$ per grid point, where $\xi=\hbar/\sqrt{m\mu}$ is the condensate healing length.
@@
with $\alpha=x,y$, $\gamma_{0}=0.1$, $w_{x}=25\xi$, $w_{y}=20\xi$ and $d=3\xi$. Furthermore, to prevent vortex reentry under the periodic boundary conditions, we use the vortex-unwinding algorithm to remove the vortex adjacent to the boundary region [33, 24]. Vortex removal is implemented every $5\tau$ during time evolution.
To investigate the long-time steady-state shedding regime, we examined the vortex shedding over a window of $2000\tau\leq t\leq 10000\tau$, thus excluding the initial acceleration stage. To break symmetry during initial shedding, we added a small amount of complex random noise to the initial wave function after the imaginary-time evolution, with uniformly distributed real and imaginary parts scaled to a relative amplitude of the order of $10^{-4}$.
@@ III.1 Vortex dipole–to–cluster transition
Figure 1: Vortex shedding in a Bose-Einstein condensate flowing past a penetrable obstacle. The condensate flows from right to left (black arrow) at a speed of $v_{0}$. The Gaussian obstacle has strength $V_{0}/\mu=0.9$ and width $\sigma/\xi=20$. (a)–(c) Superfluid vorticity distributions $\omega_{\mathrm{s}}(\mathbf{r})$ for $v_{0}/c_{s0}=0.25$ (a), $0.30$ (b), and $0.40$ (c), where $c_{s0}$ is the speed of sound in the unperturbed condensate. $\omega_{\mathrm{s}}(\mathbf{r})$ is normalized by $\omega_{\mathrm{s0}}=\rho_{0}/\tau$. As $v_{0}$ increases, shedding evolves from periodic vortex-dipole emission to same-sign vortex-cluster shedding. Dashed circles indicate the obstacle’s $1/e^{2}$ radius, centered at $x=0$. (d)–(f) Time traces of drag force $F_{\mathrm{D}}$ (solid blue) and lift force $F_{\mathrm{L}}$ (dotted red) for the same parameters used in (a)–(c). The force is normalized by $F_{\mathrm{\mu}}=\mu/\xi$. As $v_{0}$ increases, $F_{\mathrm{D}}$ grows in magnitude and becomes increasingly irregular, while $F_{\mathrm{L}}$ increases in magnitude and develops stronger oscillations. Together, these trends reflect the transition from dipole- to cluster-dominated shedding.
@@
providing a convenient measure of vortex generation, annihilation, and propagation around the obstacle. The force exerted by the obstacle is estimated as
@@
In Fig. 1, we display $\omega_{\mathrm{s}}(\mathbf{r})$ and $\mathbf{F}(t)$ for different flow velocities with $V_{0}/\mu=0.9$ and $\sigma/\xi=20$. Just above the critical velocity $v_{\mathrm{c}}\approx 0.18c_{\rm s0}$, where $c_{\rm s0}=\sqrt{\mu/m}$ is the speed of sound in the unperturbed condensate, vortex dipoles are nucleated within the penetrable obstacle via phase accumulation and phase slip [Fig. 1(a)[20]. These dipoles form nearly parallel rows as they propagate downstream. Because such dipole rows are unstable [36, 17], the downstream wake gradually reorganizes into a staggered V–shaped pattern [21]. As the velocity increases, this instability spreads upstream toward the obstacle, resulting in increasingly irregular vortex configurations in the near wake [Fig. 1(b)]. At still higher velocities, the nucleation mechanism remains predominantly pair production at the center of the obstacle, while the increased rate of pair-production raises the vortex density in the wake. The resulting high vortex density promotes vortex-vortex interactions, leading to annihilation events and the formation of Jones–Roberts solitons [37] [Fig. 1(c)]. These interactions lead to the formation of same–sign vortex clusters, which persist downstream and may eventually organize into a quasi–von Kármán street [38], despite the fact that the obstacle is penetrable (see Appendix A).
@@
The corresponding force response, shown in Figs. 1(d)1(f), reflects the same progression. In the dipole–row regime near $v_{\mathrm{c}}$, the drag force exhibits periodic oscillations with negligible lift, consistent with symmetric dipole nucleation. As the velocity increases, the drag-oscillation frequency rises, and the drag force shows larger fluctuations, in line with the growing instability in the wake. Once vortex clusters appear, the drag force becomes strongly irregular, while the lift force develops pronounced periodic oscillations associated with alternating cluster formation [Fig. 1(f)]. These results show that, as the flow velocity increases, the shedding behavior of a penetrable obstacle changes from periodic dipole–row emission to the shedding of same–sign vortex clusters. These observations are consistent with earlier numerical studies [39].
@@
We investigate the dipole–to–cluster transition further by examining the velocity dependence of the drag and lift forces. Figure 2(a) shows the time-averaged drag force $\overline{F_{\mathrm{D}}}$ and the root-mean-square of the lift force, $(\overline{F_{\mathrm{L}}^{2}})^{1/2}$ as functions of $v_{0}$. Here, the time averages are evaluated for $2000\leq t/\tau\leq 10000$ with snapshots sampled every $\Delta t=10\,\tau$. Both the drag and lift exhibit sudden increases at the threshold velocity, $v_{\mathrm{th}}\simeq 0.33\,c_{\mathrm{s0}}$, which is also associated with the onset of asymmetric vortex cluster shedding. We find that the drag force closely follows the scaling $\overline{F_{\mathrm{D}}}\sim(v_{0}-v_{\mathrm{c}})^{1/2}$ in the low-velocity dipole-emission regime [40], while it crossed over to a quadratic dependence $\overline{F_{\mathrm{D}}}\sim(v_{0}-v_{\mathrm{c}})^{2}$ in the high-velocity cluster-shedding regime [Fig. 2(a) inset].
@@
The dipole-to-cluster transition is also examined by analyzing the oscillatory behavior of the drag and lift forces through their Fourier spectra. The drag force displays a well-defined shedding frequency up to $v_{\mathrm{th}}$, consistent with the periodic dipole emission, while above $v_{\mathrm{th}}$, the spectrum becomes irregular and no dominant frequency remains. In contrast, the lift force shows a negligible spectral weight below $v_{\mathrm{th}}$ but develops a clear oscillation frequency above $v_{\mathrm{th}}$, reflecting the alternating release of vortex clusters. From the spectral peaks of the Fourier spectra of $F_{\mathrm{D}}(t)$ and $F_{\mathrm{L}}(t)$, we determine their characteristic oscillating frequencies $f_{\mathrm{D}}$ and $f_{\mathrm{L}}$, respectively. The measurement results are shown in Fig. 2(b) for the corresponding shedding regimes across the threshold velocity. In the dipole shedding regime, $f_{\mathrm{D}}$ increases almost linearly with an increase of $v_{0}$ above $v_{c}$, consistent with previous experiments [26, 28]. By contrast, in the cluster shedding regime, $f_{\mathrm{L}}$ exhibits very weak dependence on $v_{0}$. Near the threshold velocity $v_{\mathrm{th}}$, the measured frequencies suggest an approximate relation $f_{\mathrm{L}}\sim f_{\mathrm{D}}/2$, which is consistent with the onset of alternating, symmetry-broken shedding.